If sin(A-B) = ½, cos(A+B)= ½, 0B, find A and B.

Grade:10

2 Answers

Vijay Mukati
7 years ago
Assume A-B = alpha and A+B = beta.
Now, Since sin (alpha) = cos (beta) = ½. Therefore alpha = beta = pi/4
Now solve the two equations A-B = pi/4 and A+B = pi/4. and find A and B.

Thanks.
mohan
7 years ago
Hi friend,
the answer is A=pi/4
and B=0
THANKS Friend............................................................